# Pengcheng Laboratory, China

### v2

@@ 1 Introduction
    We systematically reviewed the state-of-the-art work in the intersection of software engineering and LLM, a highly discussed topic. We manually selected 123 relevant works from many articles in six databases and conducted detailed organization and categorization. (Open source, $\mathrm{https://github.com/KevinHeiwa/LLM-SE-Paper-List}$);
@@ 3.3 Data Analysis
To validate our classification, we organized the research papers from ICSE 2024  [^7] https://conf.researchr.org/track/icse-2024/icse-2024-research-track?#event-overview, a highly recognized conference in the field of software engineering, as shown in Table 3. We can observe that there were a total of 40 papers related to LLM in ICSE 2024, which can be categorized into the seven classes mentioned in our article. This indicates that the software engineering tasks we selected are currently the most prominent in LLM and SE research and encompass a significant portion of software engineering tasks. Furthermore, concerning benchmarking, datasets, and similar work, we do not categorize them as software engineering tasks. Instead, we classify them based on their research targets. For example, a benchmark for code generation would still fall under the code generation task. The definitions of different tasks are provided in Table 5.
@@ 4 LLMs in Software Engineering Tasks
In this article, we consider transformer based language models with a parameter count greater than or equal to 0.8 billion (0.8b) as Large Language Models (LLMs). We arrived at this definition based on our investigation of the majority of open-source LLM papers from ICSE 2024, as shown in Table 6. We can see that, except for OpenAI’s models, which are not open-source, we cannot determine their parameter counts. Among the open-source models, the one with the smallest parameter count is Flan-T5-small, with 80 million (80M) parameters. Therefore, we have defined LLM based on this model as the standard. Among the peer-reviewed articles we investigated, the smallest parameter count used for LLM was 80M ((Ma et al., 2024a), Flan-T5-small).
@@ 4.5 Code Evaluation
In current research on applying LLMs to code testing, a significant portion of the work focuses on test case generation. For example, empirical studies have explored the unit test generation capabilities of LLMs like ChatGPT (Yuan et al., 2023; Tang et al., 2023b), or utilizing LLMs to generate test cases (Zhao et al., 2023a; Lemieux et al., 2023; Chen et al., 2022a; Tu et al., 2023; Li et al., 2023d).
@@ 4.6 Code Management
Due to the various challenges in code management tasks, some cutting-edge works aim to leverage the power of LLMs to alleviate the complexity of the code management process, and even achieve automated code management without manual intervention (Xiao et al., 2023; Bi et al., 2023). For example, Toufique et al.(Ahmed et al., 2023) evaluates the effectiveness of LLMs in assisting engineers in managing security incidents in cloud services, while Shrivastava et al.(Shrivastava et al., 2023a) addresses the issue of LLM models struggling to understand the context present in repositories.
Additionally, some works based on LLMs have developed code management tools or frameworks to assist code managers, aiding in version management (Gao et al., 2023) and personnel training (Lanciano. et al., 2023).
@@ 5 Performance of LLM in SE Tasks
We have organized the conclusions derived from the work above, as shown in Table 7 and Table 8. We mainly screened the evaluation results from the abstract and conclusion sections of the papers.
Due to the inconsistency in evaluation methods and criteria for LLMs across different papers, we cannot directly standardize the assessment of a LLM’s performance. We convert the results from the original papers into the ”author’s confidence” of LLMs’ performance in completing SE tasks. We categorized the results into: High Confidence, Low Confidenceand Ambiguity. That is, if authors consider the performance of LLMs to be good, which is ”High Confidence”; if authors consider the performance of LLMs to be limited, which is ”Low Confidence”; If the authors cannot determine LLM’s performance, then it is ”Ambiguity”.
It is worth noting that the blue shading and the grey shading respectively indicate that, although the evaluated LLM(s) in the article did not receive a High Confidence  on the task, the article still provided a positive attitude towards its future potential; and that although LLMs is currently performing well, there are still limitations or even not good enough.
Therefore, at the current stage, LLMs still cannot fully achieve the level of professional human programmers in handling software engineering tasks, but they can serve as excellent assistants to software developers, as search tools, and as prompting tools.
@@ 7.2 External Validity
Currently, LLM technology is still evolving continuously at a very high rate. At the same time, LLM research work on SE is also developing rapidly. Although our findings cannot cover more future work. However, our work is still valuable in revealing the current direction of LLM on SE tasks. Meanwhile, our categorization of LLM on SE tasks remains applicable.
@@ 10 Data availability statements
The list of the literature surveyed in this paper is open source and is available in https://github.com/KevinHeiwa/LLM-SE-Paper-List

### v3

# Zhuhai Key Laboratory of Trusted Large Language Models

@@ 1 Introduction
    We systematically reviewed the state-of-the-art work in the intersection of software engineering and LLM, a highly discussed topic. We manually selected 123 relevant works from many articles in six databases and conducted detailed organization and categorization. (Open source, https://github.com/KevinHeiwa/LLM-SE-Paper-List);
@@ 3.3 Data Analysis
To validate our classification, we organized the research papers from ICSE 2024  [^7] https://conf.researchr.org/track/icse-2024/icse-2024-research-track?#event-overview, a highly recognized conference in the field of software engineering, as shown in Table 3. We can observe that there were a total of 40 papers related to LLM in ICSE 2024, which can be categorized into the seven classes mentioned in the paper. This indicates that the software engineering tasks we selected are currently the most prominent in LLM and SE research and encompass a significant portion of software engineering tasks. Furthermore, concerning benchmarking, datasets, and similar work, we do not categorize them as software engineering tasks. Instead, we classify them based on their research targets. For example, a benchmark for code generation would still fall under the code generation task. The definitions of different tasks are provided in Table 5.
@@ 4 LLMs in Software Engineering Tasks
In this paper, we consider models with a parameter count greater than or equal to 0.8 billion (0.8b) as Large Language Models (LLMs). We arrived at this definition based on our investigation of the majority of open-source LLM papers from ICSE 2024, as shown in Table 6. We can see that, except for OpenAI’s models, which are not open-source, we cannot determine their parameter counts. Among the open-source models, the one with the smallest parameter count is Flan-T5-small, with 80 million (80M) parameters. Therefore, we have defined LLM based on this model as the standard. Among the peer-reviewed articles we investigated, the smallest parameter count used for LLM was 80M ((Ma et al., 2024a), Flan-T5-small).
@@ 4.5 Code Evaluation
In current research on applying LLMs to code testing, a significant portion of the work focuses on test case generation. For example, empirical studies have explored the unit test generation capabilities of LLMs like ChatGPT (Yuan et al., 2023; Tang et al., 2023b) or utilizing LLMs to generate test cases (Zhao et al., 2023a; Lemieux et al., 2023; Chen et al., 2022a; Tu et al., 2023; Li et al., 2023d).
@@ 4.6 Code Management
Due to the various challenges in code management tasks, some cutting-edge works aim to leverage the power of LLMs to alleviate the complexity of the code management process, and even achieve automated code management without manual intervention (Xiao et al., 2023; Bi et al., 2023). For example, Toufique et al.(Ahmed et al., 2023) evaluates the effectiveness of LLMs in assisting engineers in managing security incidents in cloud services, while Shrivastava et al.(Shrivastava et al., 2023a) addresses the issue of LLM models struggling to understand the context present in repositories. Additionally, some works based on LLMs have developed code management tools or frameworks to assist code managers, aiding in version management (Gao et al., 2023) and personnel training (Lanciano. et al., 2023).
@@ 5 Performance of LLM in SE Tasks
We organize the conclusions derived from the work above, as shown in Table 7 and Table 8.
We mainly screen the evaluation results from the abstract and conclusion sections of the papers. Due to the inconsistency in evaluation methods and criteria for LLMs across different papers, we cannot directly standardize the assessment of a LLM’s performance. We convert the results from the original papers into the ”author’s confidence” of LLMs’ performance in completing SE tasks. We categorized the results into: High Confidence, Low Confidenceand Ambiguity. That is, if authors consider the performance of LLMs to be good, which is ”High Confidence”; if authors consider the performance of LLMs to be limited, which is ”Low Confidence”; If the authors cannot determine LLM’s performance, then it is ”Ambiguity”.
It is worth noting that the Author’s Confidence and the Author’s Confidence respectively indicate that, although the evaluated LLM(s) in the article did not receive a High Confidence  on the task, the article still provided a positive attitude towards its future potential; and that although LLMs is currently performing well, there are still limitations or even not good enough.
Therefore, at the current stage, LLMs still cannot fully achieve the level of professional human programmers in handling software engineering tasks, but they can serve as assistants to software developers, such as answering development questions and providing code examples.
@@ 7.1 Internal Validity
In Section 3, Our classification method of the SE tasks serves as an early-stage exploration to better structure this emerging field. We acknowledge that our classification may not fully encompass all aspects of software engineering tasks, and we agree that such a classification should be grounded in practice and mapped to actual research papers, rather than derived from conference categories alone. However, we believe our classification still captures a significant portion of the most important tasks in software engineering.
@@ 7.2 External Validity
Currently, LLM technology is evolving rapidly, as is LLM research in the field of software engineering. Although our findings cannot encompass future works, our work is still valuable in revealing the current direction of LLM on SE tasks. Meanwhile, our categorization of LLM in SE tasks remains applicable within the current scope of research.
@@ 10 Data availability statements
The list of the literature surveyed in this paper is open source and is available in https://github.com/KevinHeiwa/LLM-SE-Paper-List.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant 62306344, and Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515010253.